OpenStudy (anonymous):

5. What is the simplified form of –3z^2(z + 2) – 4(z^2 + 1)?

OpenStudy (callisto):

–3z^2(z + 2) – 4(z^2 + 1) = –3z^2(z) –3z^2( 2) – 4(z^2 ) -4(1) = -3z^3 -6z^2 - 4z^2 -4 = -3z^3 - 10z^2 -4

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

-3z^3 - 6z^2 - 4z^2 - 4 -3z^3 - 10z^2 -4 or u can get further help through www.teachon.tk

OpenStudy (anonymous):

thanks